OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 8x and y = 2x + 2 intersect are the solutions of the equation 8x = 2x + 2. (4 points) Part B Make tables to find the solution to 8x = 2x + 2. Take the integer values of x between -3 and 3. (4 points) Part C: How can you solve the equation 8x = 2x + 2 graphically? (2 points) @chlobohoe

OpenStudy (anonymous):

i have arrived. heehee give me a sec

OpenStudy (anonymous):

sure lol

OpenStudy (anonymous):

sorry I hate this question its supposed to be 8^x and 2^x+2

OpenStudy (anonymous):

no no its okay that literally changes nothing except the way i set up the equation. lemme say all that again hold on

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

When the two graphs of equations y=8^x and y=2^x+2 intersect, the y coordinates will have same values. That makes the right hand side expressions equal to each other. So for solving the equation 8^x=2^x+2, we can graph the equations y=8^x and y=2^x+2 and find the point which they intersect. That helps with part A AND C

OpenStudy (anonymous):

What is Part B?

OpenStudy (anonymous):

Part B basically what you have to do is make a table to find the solution to 8x = 2x + 2. What It means by "take the integer values of x between -3 and 3." is that the only x values you'll use as imputs and gonna be -3, -2, -1, 0, 1, 2, 3. okay?

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

 now all you have to do is plug in the x values to the equation 8^x=2^x+2 and you will get the y output as an answer! You'll know you did something right when you get 8 for one of the y outputs i think